# Inference-Cost-Aware Dynamic Tree Construction for Efficient Inference in Large Language Models

### v1

@@ Abstract
Therefore, we introduce a new dynamic tree decoding approach called CAST that takes into account inference costs, including factors such as GPU configurations and batch sizes, to dynamically refine the tree structure. Through comprehensive experimentation across six diverse tasks and utilizing six distinct LLMs, our methodology demonstrates remarkable results, achieving speeds up to 5.2 times faster than conventional decoding methods. Moreover, it generally outperforms existing state-of-the-art techniques from $5\%$ to $20\%$.
@@ 1 Introduction
To address this challenge, speculative decoding techniques Leviathan et al. (2023); Chen et al. (2023) have emerged. These methods swiftly generate initial tokens and validate them concurrently, thereby diminishing inference latency by generating multiple tokens in a single forward pass. While traditional speculative decoding adopts a chain-structured draft, recent progressions have introduced tree-structured drafts to boost efficiency. For example, EAGLE Li et al. (2024b) utilizes a static draft tree structure, incorporating a fixed number of candidates at each stage. However, this fixed approach overlooks the context-specific nature of token acceptance rates, contradicting the fundamental premise of speculative sampling that simpler tokens can be predicted by smaller models. Subsequently, EAGLE-2 Li et al. (2024c) and EAGLE-3 Li et al. (2025) leverage dynamic trees to further enhance performance.
@@
Our comprehensive evaluations span six distinct tasks: multi-turn conversation, code generation, mathematical reasoning, instruction following, summarization, and question answering. The datasets utilized encompass MT-bench Zheng et al. (2023), HumanEval Chen et al. (2021), GSM8K Cobbe et al. (2021), Alpaca Taori et al. (2023), CNN/Daily Mail Nallapati et al. (2016), and Natural Questions Kwiatkowski et al. (2019). We benchmark our method against state-of-the-art speculative decoding techniques: standard speculative decoding Joao Gante (2023); Leviathan et al. (2023); Chen et al. (2023), Medusa Cai et al. (2024), PLD Saxena (2023), Lookahead Fu et al. (2023), EAGLE Li et al. (2024b), EAGLE-2 Li et al. (2024c), and EAGLE-3 Li et al. (2025). Experiments are conducted across various LLM series with different batch sizes, including Vicuna, LLaMA3, Qwen2, and distilled DeepSeek-R1. Our method consistently surpasses all baseline approaches, achieving speedups of up to 5.2x and typically delivering speed enhancements ranging from $5\%$ to $20\%$ compared to the previous state-of-the-art method.
@@ 2 Related Works
Subsequent methods improved draft efficiency and base-model alignment. SpecInfer Miao et al. (2023) used draft-model ensembles and tree-mask attention. Medusa Cai et al. (2024) leveraged MLPs on internal states to predict multiple tokens. EAGLE Li et al. (2024b) expanded tree proposals for higher acceptance. Draft-and-Verify frameworks Zhang et al. (2023); Hooper et al. (2023); Yang et al. (2023); Monea et al. (2023); Li et al. (2024a); Yi et al. (2024); Liu et al. (2024); Sun et al. (2024); Elhoushi et al. (2024); Svirschevski et al. (2024) introduced early exits and partial model reuse, partitioning the LLM into fast generators and verifiers.
More recently, dynamic draft trees emerged. GLIDE and CAPE Du et al. (2024) added fallback branches for uncertain cases but limited expansion. EAGLE-2 Li et al. (2024c) removed such constraints for fully adaptive growth, while EAGLE-3 Li et al. (2025) further relaxed training restrictions, yielding more effective speculative decoding.
@@ 3.2 EAGLE
The previously discussed speculative decoding method predicts the tokens in an autoregressive chain and verifies them sequentially. It has the disadvantage of once rejecting a token, all its subsequent tokens will also be discarded. EAGLE Li et al. (2024b) improves speculative decoding by constructing a tree-structured draft and performing parallel verification, the tree structure makes the rejection process still retain some information by leaving the tokens in the rejected token’s sibling subtree un-discarded.
Unlike EAGLE Li et al. (2024b), which uses a predefined static tree, EAGLE-2 (Li et al., 2024c) and EAGLE-3 (Li et al., 2025) improve speculative decoding by dynamically constructing a tree-structured draft using the confidence score. The dynamic structure makes the inference much more data-dependent and performs much better. We will then briefly discuss some of the details.
@@ 4.1 Dynamic Expansion Stage: Breadth and Depth Pruning
Empirically, the acceptance rate of a node $u$ is strongly correlated with its confidence score $v(u)$ (Du et al., 2024; Li et al., 2024c), which we use as a proxy for acceptance probability. Drawing inspiration from utility theory in economics, we frame node selection as a utility maximization problem.
@@ 5 Experiments
Following prior works, we perform experiments on a variety of models across diverse sizes, including Vicuna-13B-v1.3 (V 13B) and Vicuna-33B-v1.3(V 33B)Chiang et al. (2023),Llama-3.1-8B-Instruct(L31 8B) and Llama-3.3-70B-Instruct(L33 70B) Meta (2024), DeepSeek-R1-Distill-Llama-8B(DSL 8B), Qwen2-7B-Instruct(Q2 7B). Following the standard benchmark in this area, we conducted extensive evaluations across six different text generation tasks to show the applicability of our method under diverse scenarios, includng multi-turn conversation, code generation, mathematical reasoning, instruction following, summarization, and question answering, we used the MT-bench Zheng et al. (2023), HumanEval Chen et al. (2021), GSM8K Cobbe et al. (2021), Alpaca Taori et al. (2023), CNN/Daily Mail Nallapati et al. (2016), and Natural Questions Kwiatkowski et al. (2019) datasets, respectively. In line with common practices in the community, we employed the same initial model weights for all tasks without any modifications. We will use the vanilla autoregressive decoding as the baseline for comparison, with a speedup ratio of 1.00x. Our method will be compared with the most state-of-the-art methods in speculative decoding and we will use their default hyperparameters, including standard speculative decoding (SpD) Joao Gante (2023); Leviathan et al. (2023); Chen et al. (2023), Medusa Cai et al. (2024), PLD Saxena (2023), Lookahead Fu et al. (2023), EAGLE Li et al. (2024b), EAGLE-2 Li et al. (2024c), and EAGLE-3 Li et al. (2025). Under non-greedy setting, methods like Medusa relax acceptance condition, so we will not compare with method like this. Given that speedup ratios are hardware-dependent, we ensured fairness by testing all methods on identical devices, which are the Nvidia A800 GPUs. All experiments run relatively fast, usually less than one hour, even for large datasets, as only inference is performed. The experiments on more GPU types can be found in Appendix.

### v2

@@ Abstract
Therefore, we introduce a new dynamic tree decoding approach called CAST that takes into account inference costs, including factors such as GPU configurations and batch sizes, to dynamically refine the tree structure. Through comprehensive experimentation across six diverse tasks and utilizing six distinct LLMs, our methodology demonstrates remarkable results, achieving speeds up to 5.2 times faster than conventional decoding methods. Moreover, it generally outperforms existing state-of-the-art techniques from $5\%$ to $20\%$. The code is available at https://github.com/EAGLE-Research/sglang-eagle4.
@@ 1 Introduction
To address this challenge, speculative decoding techniques Leviathan et al. (2023); Chen et al. (2023) have emerged. These methods swiftly generate initial tokens and validate them concurrently, thereby diminishing inference latency by generating multiple tokens in a single forward pass. While traditional speculative decoding adopts a chain-structured draft, recent progressions have introduced tree-structured drafts to boost efficiency. For example, EAGLE Li et al. (2024c) utilizes a static draft tree structure, incorporating a fixed number of candidates at each stage. However, this fixed approach overlooks the context-specific nature of token acceptance rates, contradicting the fundamental premise of speculative sampling that simpler tokens can be predicted by smaller models. Subsequently, EAGLE-2 Li et al. (2024b) and EAGLE-3 Li et al. (2025) leverage dynamic trees to further enhance performance.
@@
Our comprehensive evaluations span six distinct tasks: multi-turn conversation, code generation, mathematical reasoning, instruction following, summarization, and question answering. The datasets utilized encompass MT-bench Zheng et al. (2023), HumanEval Chen et al. (2021), GSM8K Cobbe et al. (2021), Alpaca Taori et al. (2023), CNN/Daily Mail Nallapati et al. (2016), and Natural Questions Kwiatkowski et al. (2019). We benchmark our method against state-of-the-art speculative decoding techniques: standard speculative decoding Joao Gante (2023); Leviathan et al. (2023); Chen et al. (2023), Medusa Cai et al. (2024), PLD Saxena (2023), Lookahead Fu et al. (2023), EAGLE Li et al. (2024c), EAGLE-2 Li et al. (2024b), and EAGLE-3 Li et al. (2025). Experiments are conducted across various LLM series with different batch sizes, including Vicuna, LLaMA3, Qwen2, and distilled DeepSeek-R1. Our method consistently surpasses all baseline approaches, achieving speedups of up to 5.2x and typically delivering speed enhancements ranging from $5\%$ to $20\%$ compared to the previous state-of-the-art method.
@@ 2 Related Works
Subsequent methods improved draft efficiency and base-model alignment. SpecInfer Miao et al. (2023) used draft-model ensembles and tree-mask attention. Medusa Cai et al. (2024) leveraged MLPs on internal states to predict multiple tokens. EAGLE Li et al. (2024c) expanded tree proposals for higher acceptance. Draft-and-Verify frameworks Zhang et al. (2023); Hooper et al. (2023); Yang et al. (2023); Monea et al. (2023); Li et al. (2024a); Yi et al. (2024); Liu et al. (2024); Sun et al. (2024); Elhoushi et al. (2024); Svirschevski et al. (2024) introduced early exits and partial model reuse, partitioning the LLM into fast generators and verifiers.
More recently, dynamic draft trees emerged. GLIDE and CAPE Du et al. (2024) added fallback branches for uncertain cases but limited expansion. EAGLE-2 Li et al. (2024b) removed such constraints for fully adaptive growth, while EAGLE-3 Li et al. (2025) further relaxed training restrictions, yielding more effective speculative decoding.
@@ 3.2 EAGLE
The previously discussed speculative decoding method predicts the tokens in an autoregressive chain and verifies them sequentially. It has the disadvantage of once rejecting a token, all its subsequent tokens will also be discarded. EAGLE Li et al. (2024c) improves speculative decoding by constructing a tree-structured draft and performing parallel verification, the tree structure makes the rejection process still retain some information by leaving the tokens in the rejected token’s sibling subtree un-discarded.
Unlike EAGLE Li et al. (2024c), which uses a predefined static tree, EAGLE-2 (Li et al., 2024b) and EAGLE-3 (Li et al., 2025) improve speculative decoding by dynamically constructing a tree-structured draft using the confidence score. The dynamic structure makes the inference much more data-dependent and performs much better. We will then briefly discuss some of the details.
@@ 4.1 Dynamic Expansion Stage: Breadth and Depth Pruning
Empirically, the acceptance rate of a node $u$ is strongly correlated with its confidence score $v(u)$ (Du et al., 2024; Li et al., 2024b), which we use as a proxy for acceptance probability. Drawing inspiration from utility theory in economics, we frame node selection as a utility maximization problem.
@@ 5 Experiments
Following prior works, we perform experiments on a variety of models across diverse sizes, including Vicuna-13B-v1.3 (V 13B) and Vicuna-33B-v1.3(V 33B)Chiang et al. (2023),Llama-3.1-8B-Instruct(L31 8B) and Llama-3.3-70B-Instruct(L33 70B) Meta (2024), DeepSeek-R1-Distill-Llama-8B(DSL 8B), Qwen2-7B-Instruct(Q2 7B). Following the standard benchmark in this area, we conducted extensive evaluations across six different text generation tasks to show the applicability of our method under diverse scenarios, includng multi-turn conversation, code generation, mathematical reasoning, instruction following, summarization, and question answering, we used the MT-bench Zheng et al. (2023), HumanEval Chen et al. (2021), GSM8K Cobbe et al. (2021), Alpaca Taori et al. (2023), CNN/Daily Mail Nallapati et al. (2016), and Natural Questions Kwiatkowski et al. (2019) datasets, respectively. In line with common practices in the community, we employed the same initial model weights for all tasks without any modifications. We will use the vanilla autoregressive decoding as the baseline for comparison, with a speedup ratio of 1.00x. Our method will be compared with the most state-of-the-art methods in speculative decoding and we will use their default hyperparameters, including standard speculative decoding (SpD) Joao Gante (2023); Leviathan et al. (2023); Chen et al. (2023), Medusa Cai et al. (2024), PLD Saxena (2023), Lookahead Fu et al. (2023), EAGLE Li et al. (2024c), EAGLE-2 Li et al. (2024b), and EAGLE-3 Li et al. (2025). Under non-greedy setting, methods like Medusa relax acceptance condition, so we will not compare with method like this. Given that speedup ratios are hardware-dependent, we ensured fairness by testing all methods on identical devices, which are the Nvidia A800 GPUs. All experiments run relatively fast, usually less than one hour, even for large datasets, as only inference is performed. The experiments on more GPU types can be found in Appendix.
@@ 6 Conclusion
Acknowledgments
This work was supported by the Science and Technology Major Project of Yunnan Province (202302AF080006), the National Key R&D Program of China (2022YFB2703500).